Determine whether the lines x-5y=15 and -5x - y = 2 = 0 are parallel, perpendicular, or neither.
Answer by stanbon(75887)

Determine whether the lines x-5y=15 and -5x - y = 2 are parallel, perpendicular, or neither.
---
Compare the slopes:
y = (1/5)x-3
y = -5x -2
-----
(1/5)(-5) = 1 so the lines are perpendicular.
================
Cheers,
Stan H.
==================